OpenStudy (anonymous):

Which of the following is true? Answer x=45 x>45 x<45 None of these.

OpenStudy (anonymous):

\[\large \tt{What's\space the\space\ question?}\]

OpenStudy (anonymous):

x>45

OpenStudy (anonymous):

tan(x)=21/20 ->just over 1 tan (45)=1 and tan=sin/cos because tan >1 means sin>cos => over 45

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

ok thanks